OpenStudy (anonymous):

Find each unit rate. A store sells 3 video tapes for $4.99 & Twenty-five students use 120 sheets of paper. PLEASE HELP

OpenStudy (anonymous):

make fractions with numbers & units separate and then just simplify the numerical portions: You can read the first example as 4.99 dollars per 3 video tapes, which is confusing. \[\frac{4.99}{3}\frac{$}{videoTape}\] \[1.66\frac{$}{videoTape}\] which can now be read as 1.66 dollars per videoTape